OpenStudy (daniellelovee):

I anyone here good with chemistry? Please I need help but no one seems to reply on the chemistry section :(

OpenStudy (one098):

I'll help.

OpenStudy (daniellelovee):

THANK YOU! ok so the thing is that I have to do this experiment where I put a ballon into an ice bucket and measure its temperature but the lab says 1. From Part 2 of the lab activity, plot a graph of volume vs. temperature (in kelvins) with the two data points that resulted from your experiment. Draw a straight line connecting the two points, thus assuming a linear relationship between volume and temperature. The x-intercept corresponds to where the volume would be zero and the temperature can be no lower, defined as absolute zero. Absolute zero is 0 K. Compare your results to those expected. How close to absolute zero was your intercept? Why might your value be different from absolute zero? I already have my measurements but still the 0K part is confusing me

OpenStudy (one098):

Oh hey, would you mind re-posting this ^^ in the chemistry section? Just to avoid trouble. Thank you!!

OpenStudy (daniellelovee):

this are my results balloon before: 15 cm, C=52.25pie balloon after: 14 cm, C=49pie room: 19.4 Cº 292.4 K water: 5 Cº 278 K V = (4/3)pi(r^3) C=2pir R=C/2pi R for before=8.32 R for after=7.80 Volume for before: 2411.228472 cm^3 volume for after: 1986.79104 cm^3

OpenStudy (daniellelovee):

oh yeah sorry